# On the kissing number of the cross-polytope

### v1

@@ 1 Introduction
The translative kissing number problem asks for the maximum number of non-overlapping translates of a given convex body that can touch a central one. The kissing number of the $n$-dimensional $\ell^{2}$-ball $B_{n}=\{x\in\mathbb{R}^{n}:\sum_{i=1}^{n}x_{i}^{2}\leq 1\}$ is known in dimensions 11114444, $8$ and $24$ [1, 2, 3, 4], and the asymptotic bounds $2^{0.2075{(1+o(1))n}}\leq\kappa_{T}(B_{n})\leq 2^{0.401(1+o(1))n}$ follow from the works of Wyner [5], and Kabatiansky and Levenshtein [6] and Levenshtein [3]. The kissing number problem for convex bodies different from $B_{n}$ are equally interesting and important. See [7] for a survey on kissing numbers of convex bodies and related problems.
@@
This seems to be the first improvement on Hadwiger’s bound for $\ell^{p}$-balls different from $B_{n}$ (note that $\kappa_{T}(Q_{n})=3^{n}-1$, where $Q_{n}$ is the $n$-cube). The bound $\kappa_{L}(C)\leq 2(2^{n}-1)$ for strictly convex $C$ is a classical bound by Minkowski [11], see also [12, Theorem 30.2]. While $K_{n}$ is not strictly convex, it is possible to obtain an equally strong inequality for $\kappa_{L}(K_{n})$, up to a constant factor. See Section 3.
@@ 2 Preliminaries
The maximum size of a kissing configuration is denoted $\kappa_{T}(C)$ and sometimes called the Hadwiger number of $C$. The maximum size of a kissing configuration of $C$ consisting of the shortest non-zero vectors (with respect to the norm $\lVert\cdot\rVert_{C}$) of a lattice is denoted $\kappa_{L}(C)$. Recall that a lattice $\Lambda$ is the $\mathbb{Z}$-span of a basis of $\mathbb{R}^{n}$. Clearly, $\kappa_{L}(C)\leq\kappa_{T}(C)$. The condition that the shortest vectors of a lattice $\Lambda$ form a kissing configuration is equivalent to the condition that $\min_{x\in\Lambda\setminus\{0\}}\lVert x\rVert_{C}=1$, and the condition that the kissing configuration is as large as possible is equivalent to the set $\mathcal{M}_{C}(\Lambda)=\{x\in\Lambda:\lVert x\rVert_{C}=1\}$ having maximal cardinality.
We call two lattices $\Lambda_{1}$ and $\Lambda_{2}$ $C$-equivalent, if $\Lambda_{1}=\sigma(\Lambda_{2})$ for some $\sigma\in\mathcal{G}_{C}$. Note that $|\mathcal{M}_{C}(\Lambda)|$ depends only on the $C$-equivalence class of $\Lambda$. Define the support of a vector $x=(x_{1},\dots,x_{n})\in\mathbb{R}^{n}$ to be the set $\text{supp}(x)=\{i:x_{i}\neq 0\}$. We say that $i$ is a max-coordinate of $x$ if $|x_{i}|=\lVert x\rVert_{\infty}$. Finally, define the set $D_{n}^{+}\subset\mathbb{R}^{n}$ to be the union of $D_{n}$ and its translate $D_{n}+(\frac{1}{2},\dots,\frac{1}{2})$, where $D_{n}=\{x\in\mathbb{Z}^{n}:\sum_{i=1}^{n}x_{i}\equiv 0\pmod{2}\}$. It is easily seen that $D_{n}^{+}$ is a lattice if and only if $n$ is even. For instance, $D_{8}^{+}$ is isometric to the $E_{8}$ lattice, while $D_{4}^{+}$ is a scaled Hadamard rotation of the integer lattice $\mathbb{Z}^{4}$.
@@ 3 The lattice kissing number of $K_{n}$
Let $L\subset\mathbb{R}^{4}$ be a lattice achieving $\kappa_{L}(K_{4})$, and let $\mathcal{M}$ be a set containing exactly one of the $\pm$ pairs of vectors in $\mathcal{M}_{K_{4}}(L)$. Suppose that $|\mathcal{M}|\geq 20$ and consider the image of $\mathcal{M}$ in $L/2L$. Note that if $x,y\in\mathcal{M}$ are such that $x\equiv y\pmod{2L}$, then $\frac{x\pm y}{2}\in\mathcal{M}_{K_{4}}(L)$ and the triangle inequality implies that $\text{supp}(x)\cap\text{supp}(y)=\emptyset$, and in particular $|\text{supp}(x)|\leq 2$ or $|\text{supp}(y)|\leq 2$.
@@
Since there are $152^{4}-1=$ 15 non-zero classes in $L/2L$, the following is true: for some $0\leq k\leq 4$, $\mathcal{M}$ contains $k$ vectors $x$ with $|\text{supp}(x)|=1$, and $5-k$ pairs $\{x_{1}^{i},x_{2}^{i}\}_{i=1}^{5-k}$ of vectors such that $x_{1}^{i}\equiv x_{2}^{i}\pmod{2L}$ and $|\text{supp}(x_{1}^{i})|=|\text{supp}(x_{2}^{i})|=2$ for all $i=1,\dots,5-k$. Note that there must exist at least one pair $\{x_{1}^{i_{1}},x_{2}^{i_{1}}\}$ satisfying $\lVert x_{1}^{i_{1}}\rVert_{\infty}=\lVert x_{2}^{i_{1}}\rVert_{\infty}=\frac{1}{2}$. Indeed, if $x,y\in\mathcal{M}$ share a max-coordinate, then $\lVert x\rVert_{\infty}\leq\frac{1}{2}$ or $\lVert y\rVert_{\infty}\leq\frac{1}{2}$, and every $x\in\mathcal{M}$ has $|x_{i}|\leq\frac{1}{2}$ for all $i\in\cup_{x\in\mathcal{M}:|\text{supp}(x)|=1}\text{supp}(x)$. If $k\leq 2$, then in addition to the pair $\{x_{1}^{i_{1}},x_{2}^{i_{1}}\}$ there are two vectors $x_{2},x_{3}\in\mathcal{M}$ such that $\lVert x_{2}\rVert_{\infty}=\lVert x_{3}\rVert_{\infty}=\frac{1}{2}$ and $|\text{supp}(x_{2})|=|\text{supp}(x_{3})|=2$.
Thus, replacing $L$ by a $K_{4}$-equivalent lattice, we may assume that $L$ contains the lattice $L^{\prime}$ generated by $\mathbb{Z}^{4}$ and the vectors $(\frac{1}{2},\frac{1}{2},0,0)$, $(0,0,\frac{1}{2},\frac{1}{2})$ and their midpoint $(\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4})$. Note that $\frac{|\mathcal{M}_{K_{4}}(L^{\prime})|}{2}=10$; thus, since $L^{\prime}$ has covering radius $1$ by the next lemma, $L$ must contain in addition to $L^{\prime}$ some of its deep holes. Thus, by the next lemma, $L$ contains $\frac{1}{2}D_{4}^{+}$. The claim follows by noting that $\frac{1}{2}D_{4}^{+}$ has $\ell^{1}$-norm covering radius $<1$. ∎
In the next lemma, $\mathcal{H}_{2}$ is the planar lattice generated by $(1,0)$ and $(\frac{1}{2},\frac{1}{2})$. By deep holes of a discrete set $S\subset\mathbb{R}^{n}$, we refer to the points $y\in\mathbb{R}^{n}$ achieving the supremum $\sup_{y\in\mathbb{R}^{n}}\inf_{x\in S}\lVert x-y\rVert_{1}$, which we call the ($\ell^{1}$-norm) covering radius of $S$.
@@
The set of deep holes of the lattice $L^{\prime}=(\mathcal{H}_{2}\oplus\mathcal{H}_{2})\cup((\mathcal{H}_{2}\oplus\mathcal{H}_{2})+(\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4}))$ is the translate $(\frac{1}{2},0,\frac{1}{2},0)+\mathcal{H}_{2}\oplus\mathcal{H}_{2}$.
@@
Note that the set of deep holes of $\mathcal{H}_{2}$ is the union of translates by vectors in $\mathcal{H}_{2}$ of the line segments {±$\{\pm(x,\frac{1}{2}-x):0\leq x\leq\frac{1}{2}\}$. Thus, the deep holes of $\mathcal{H}_{2}\oplus\mathcal{H}_{2}$ are of the form $x=(\pm(x_{1},\frac{1}{2}-x_{1}),\pm(x_{2},\frac{1}{2}-x_{2}))+v$, where $0\leq x_{1},x_{2}\leq\frac{1}{2}$ and $v\in\mathcal{H}_{2}\oplus\mathcal{H}_{2}$. In particular, $\mathcal{H}_{2}\oplus\mathcal{H}_{2}$ has covering radius $1$, and since the same is true for $L^{\prime}$, the deep holes of $L^{\prime}$ are among those of $\mathcal{H}_{2}\oplus\mathcal{H}_{2}$. Since
$\inf_{0\leq x_{1},x_{2}\leq\frac{1}{2}}\lVert(x_{1},\frac{1}{2}-x_{1},x_{2},\frac{1}{2}-x_{2})-(\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4})\rVert_{1}=1$
is attained precisely for $x_{1},x_{2}\in\{0,\frac{1}{2}\}$, the lemma follows.∎

### v2

@@ 1 Introduction
The translative kissing number problem asks for the maximum number of non-overlapping translates of a given convex body that can touch a central one. The kissing number of the $n$-dimensional $\ell^{2}$-ball $B_{n}=\{x\in\mathbb{R}^{n}:\sum_{i=1}^{n}x_{i}^{2}\leq 1\}$ is known in dimensions 11114444, $8$ and $24$ [1, 2, 3, 4], and the asymptotic bounds $2^{0.2075{(1+o(1))n}}\leq\kappa_{T}(B_{n})\leq 2^{0.401(1+o(1))n}$ follow from the works of Wyner [5], Kabatiansky and Levenshtein [6] and Levenshtein [3]. The kissing number problem for convex bodies different from $B_{n}$ are equally interesting and important. See [7] for a survey on kissing numbers of convex bodies and related problems.
@@
This seems to be the first improvement on Hadwiger’s bound for $\ell^{p}$-balls different from $B_{n}$ (note that $\kappa_{T}(Q_{n})=3^{n}-1$, where $Q_{n}$ is the $n$-cube). The lattice kissing number bound $\kappa_{L}(C)\leq 2(2^{n}-1)$ for strictly convex $C$ is a classical bound by Minkowski [11], see also [12, Theorem 30.2]. While $K_{n}$ is not strictly convex, it is possible to obtain an equally strong inequality for $\kappa_{L}(K_{n})$, up to a constant factor. See Section 3.
@@ 2 Preliminaries
The maximum size of a kissing configuration is denoted $\kappa_{T}(C)$ and sometimes called the Hadwiger number of $C$. The maximum size of a kissing configuration of $C$ consisting of the shortest non-zero vectors (with respect to the norm $\lVert\cdot\rVert_{C}$) of a lattice is denoted $\kappa_{L}(C)$. Recall that a lattice $\Lambda$ is the $\mathbb{Z}$-span of a basis of $\mathbb{R}^{n}$. Clearly, $\kappa_{L}(C)\leq\kappa_{T}(C)$. The condition that the shortest non-zero vectors of a lattice $\Lambda$ form a kissing configuration is equivalent to the condition that $\min_{x\in\Lambda\setminus\{0\}}\lVert x\rVert_{C}=1$, and the condition that the kissing configuration is as large as possible is equivalent to the set $\mathcal{M}_{C}(\Lambda)=\{x\in\Lambda:\lVert x\rVert_{C}=1\}$ having maximal cardinality.
We call two lattices $\Lambda_{1}$ and $\Lambda_{2}$ $C$-equivalent, if $\Lambda_{1}=\sigma(\Lambda_{2})$ for some $\sigma\in\mathcal{G}_{C}$. Note that $|\mathcal{M}_{C}(\Lambda)|$ depends only on the $C$-equivalence class of $\Lambda$. Define the support of a vector $x=(x_{1},\dots,x_{n})\in\mathbb{R}^{n}$ to be the set $\text{supp}(x)=\{i:x_{i}\neq 0\}$. We say that $i$ is a max-coordinate of $x$ if $|x_{i}|=\lVert x\rVert_{\infty}$. We define the set $D_{n}^{+}\subset\mathbb{R}^{n}$ to be the union of $D_{n}$ and its translate $D_{n}+(\frac{1}{2},\dots,\frac{1}{2})$, where $D_{n}=\{x\in\mathbb{Z}^{n}:\sum_{i=1}^{n}x_{i}\equiv 0\pmod{2}\}$. It is easily seen that $D_{n}^{+}$ is a lattice if and only if $n$ is even. For instance, $D_{8}^{+}$ is isometric to the $E_{8}$ lattice, while $D_{4}^{+}$ is a scaled Hadamard rotation of the integer lattice $\mathbb{Z}^{4}$. We define the ($\ell^{1}$-norm) covering radius of a discrete set $S\subset\mathbb{R}^{n}$ to be the quantity $\rho(S)=\sup_{y\in\mathbb{R}^{n}}\inf_{x\in S}\lVert x-y\rVert_{1}$. We call the points $y\in\mathbb{R}^{n}$ achieving the supremum deep holes of $S$.
@@ 3 The lattice kissing number of $K_{n}$
Let $L\subset\mathbb{R}^{4}$ be a lattice achieving $\kappa_{L}(K_{4})$, and let $\mathcal{M}$ be a set containing exactly one of each $\pm$ pairs of vectors in $\mathcal{M}_{K_{4}}(L)$. Suppose that $|\mathcal{M}|\geq 20$. Note that if $x,y\in\mathcal{M}$ are distinct vectors such that $x\equiv y\pmod{2L}$, then $\frac{x\pm y}{2}\in\mathcal{M}_{K_{4}}(L)$ and the triangle inequality implies that $\text{supp}(x)\cap\text{supp}(y)=\emptyset$, and in particular $|\text{supp}(x)|\leq 2$ or $|\text{supp}(y)|\leq 2$.
@@
Since there are $152^{4}-1=$ 15 non-zero classes in $L/2L$, the following is true: $\mathcal{M}$ contains $k\in\{0,1,2,3,4\}$ vectors whose support has size one and $5-k$ pairs $\{x_{1}^{i},x_{2}^{i}\}_{i=1}^{5-k}$ of vectors such that $x_{1}^{i}\equiv x_{2}^{i}\pmod{2L}$ and $|\text{supp}(x_{1}^{i})|=|\text{supp}(x_{2}^{i})|=2$ for all $i=1,\dots,5-k$. Note that there must exist at least one pair $\{x_{1}^{i_{1}},x_{2}^{i_{1}}\}$ satisfying $\lVert x_{1}^{i_{1}}\rVert_{\infty}=\lVert x_{2}^{i_{1}}\rVert_{\infty}=\frac{1}{2}$. Indeed, if $x,y\in\mathcal{M}$ share a max-coordinate, then $\lVert x\rVert_{\infty}\leq\frac{1}{2}$ or $\lVert y\rVert_{\infty}\leq\frac{1}{2}$, and every $x\in\mathcal{M}$ with $|\text{supp}(x)|>1$ has $|x_{i}|\leq\frac{1}{2}$ for all $i\in\cup_{x\in\mathcal{M}:|\text{supp}(x)|=1}\text{supp}(x)$. If $k\leq 2$, then in addition to the pair $\{x_{1}^{i_{1}},x_{2}^{i_{1}}\}$ there are two vectors $x_{3},x_{4}\in\mathcal{M}$ such that $\lVert x_{3}\rVert_{\infty}=\lVert x_{4}\rVert_{\infty}=\frac{1}{2}$ and $|\text{supp}(x_{3})|=|\text{supp}(x_{4})|=2$. Thus, replacing $L$ by a $K_{4}$-equivalent lattice, we may assume that $L$ contains the lattice $L^{\prime}$ generated by $\mathbb{Z}^{4}$, the vectors $(\frac{1}{2},\frac{1}{2},0,0)$, $(0,0,\frac{1}{2},\frac{1}{2})$ and their midpoint $(\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4})$.
Since $|\mathcal{M}_{K_{4}}(L^{\prime})|=20$ and $\rho(L^{\prime})\leq 1$ by the next lemma, $L$ contains in addition to $L^{\prime}$ at least one of its deep holes. By the next lemma, $L\setminus L^{\prime}$ contains a vector of one of the following shapes.
$v_{1}$ $=(x_{1},\frac{1}{2}-x_{1},x_{2},\frac{1}{2}-x_{2}),\quad 0\leq x_{1},x_{2}\leq\frac{1}{2},$
$v_{2}$ $=(x_{1},x_{1}-\frac{1}{2},x_{2},\frac{1}{2}-x_{2}),\quad 0\leq x_{1},x_{2}\leq\frac{1}{2},$
$v_{2}^{\prime}$ $=(x_{1},\frac{1}{2}-x_{1},x_{2},x_{2}-\frac{1}{2}),\quad 0\leq x_{1},x_{2}\leq\frac{1}{2},$
$v_{3}$ $=(x_{1},x_{1}-\frac{1}{2},x_{2},x_{2}-\frac{1}{2}),\quad 0\leq x_{1},x_{2}\leq\frac{1}{2}.$
Suppose $v_{1}\in L$. Then since
$\lVert(x_{1},\frac{1}{2}-x_{1},x_{2},\frac{1}{2}-x_{2})-(\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4})\rVert_{1}\leq 1$
for all $0\leq x_{1},x_{2}\leq\frac{1}{2}$ with equality if and only if $x_{1},x_{2}\in\{0,\frac{1}{2}\}$, $L$ contains $(\frac{1}{2},0,\frac{1}{2},0)$ and so $L$ contains $\frac{1}{2}D_{4}^{+}$.
Suppose $v_{2}\in L$. Then
$\lVert(x_{1},x_{1}-\frac{1}{2},x_{2},\frac{1}{2}-x_{2})-(\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4})\rVert_{1}\leq 1$
for all $\frac{1}{4}\leq x_{1}\leq\frac{1}{2}$, $0\leq x_{2}\leq\frac{1}{2}$ with equality if and only if $x_{2}\in\{0,\frac{1}{2}\}$. Similarly,
$\lVert(x_{1},x_{1}-\frac{1}{2},x_{2},\frac{1}{2}-x_{2})-(\frac{1}{4},-\frac{3}{4},\frac{1}{4},\frac{1}{4})\rVert_{1}\leq 1$
for all $0\leq x_{1}\leq\frac{1}{4}$, $0\leq x_{2}\leq\frac{1}{2}$ with equality if and only if $x_{2}\in\{0,\frac{1}{2}\}$. Thus, $L$ contains $(x_{1},x_{1}-\frac{1}{2},0,\frac{1}{2})$ or $(x_{1},x_{1}-\frac{1}{2},\frac{1}{2},0)$ so it contains $(2x_{1}-\frac{1}{2},2x_{1}-\frac{1}{2},0,0)$. Since $-\frac{1}{2}\leq 2x_{1}-\frac{1}{2}\leq\frac{1}{2}$, we must have $x_{1}\in\{0,\frac{1}{4},\frac{1}{2}\}$. Thus, $L$ contains $(\frac{1}{2},0,\frac{1}{2},0)$ or $(\frac{1}{4},-\frac{1}{4},0,\frac{1}{2})$ so it contains $\frac{1}{2}D_{4}^{+}$ or the lattice $L_{0}$ spanned by $L^{\prime}$ and $(\frac{1}{4},-\frac{1}{4},0,\frac{1}{2})$. The assumption $v_{2}^{\prime}\in L$ leads to the same conclusion.
Now suppose $v_{3}\in L$. Then $2v_{3}+(0,1,0,1)=(2x_{1},2x_{1},2x_{2},2x_{2})\in L$. It is easy to see that this implies $(2x_{1},2x_{1},2x_{2},2x_{2})\in L^{\prime}$. Then the only possibilities are $x_{1},x_{2}\in\{0,\frac{1}{4},\frac{1}{2}\}$ or $x_{1},x_{2}\in\{\frac{1}{8},\frac{3}{8}\}$. In the former case, $L$ contains $\frac{1}{2}D_{4}^{+}$ or $L_{0}$. In the latter case, $L$ contains the lattice $L_{1}$ spanned by $L^{\prime}$ and $(\frac{1}{8},-\frac{3}{8},\frac{1}{8},-\frac{3}{8})$ or the lattice $L_{1}^{\prime}$ spanned by $L^{\prime}$ and $(\frac{1}{8},-\frac{3}{8},\frac{3}{8},-\frac{1}{8})$. Note that $L_{1}$ is $K_{4}$-equivalent to $L_{1}^{\prime}$.
In conclusion, $L$ contains $\frac{1}{2}D_{4}^{+}$, $L_{0}$, $L_{1}$ or $L_{1}^{\prime}$. Suppose that $L$ strictly contains one of these lattices. Then by the reasoning above, $L$ contains at least two of these lattices. This however implies that $\min_{x\in L\setminus\{0\}}\lVert x\rVert_{1}<1$, a contradiction. In conclusion, $L\in\{\frac{1}{2}D_{4}^{+},L_{0},L_{1},L_{1}^{\prime}\}$. We check that
$|\mathcal{M}_{K_{4}}(\frac{1}{2}D_{4}^{+})|=40,\quad|\mathcal{M}_{K_{4}}(L_{0})|=36,\quad|\mathcal{M}_{K_{4}}(L_{1})|=|\mathcal{M}_{K_{4}}(L_{1}^{\prime})|=28,$
and conclude that $L=\frac{1}{2}D_{4}^{+}$.
@@
Let $\mathcal{H}_{2}$ be the planar lattice generated by $(1,0)$ and $(\frac{1}{2},\frac{1}{2})$. Then the set of deep holes of the lattice $\mathcal{H}_{2}\oplus\mathcal{H}_{2}\subset\mathbb{R}^{4}$ is the set
$\left\{(x_{1},\pm(\frac{1}{2}-x_{1}),x_{2},\pm(\frac{1}{2}-x_{2})):0\leq x_{1},x_{2}\leq\frac{1}{2}\right\}+\mathcal{H}_{2}\oplus\mathcal{H}_{2}.$
In particular, $\rho(\mathcal{H}_{2}\oplus\mathcal{H}_{2})=1$.
@@
The deep holes of a direct sum $L_{1}\oplus L_{2}$ of lattices $L_{1}$, $L_{2}$ are given by $(x,y)$, where $x$ is a deep hole of $L_{1}$ and $y$ is a deep hole of $L_{2}$. Thus, it suffices to show that the deep holes of $\mathcal{H}_{2}$ consist of the translates by vectors in $\mathcal{H}_{2}$ of the line segments $\{(x,\pm(\frac{1}{2}-x)):0\leq x\leq\frac{1}{2}\}$. This is clear. ∎